XYZ stock has a share price of $125 today. All rates of interest are 5% per year with continuous compounding. Finally, XYZ is scheduled to pay the following dividends per share over the next year: a dividend of $3.0 per share in three months and a dividend of $3.0 per share in six months. Derive today’s forward price of the stock for delivery in nine months.
Answer #1

PV of price of future payments today =Share Price -Dividend 1*e^(-r*t)- Dividend 2*e^(-rt)
=125-3*EXP(-5%*3/12)-3*(-5%*6/12) =122.1123

Today's Forward price =PV of price of future payments today*EXP(5%*9/12) =122.1123*EXP(5%*9/12) =126.78